# Eaman Jahani

### v4

@@ Results
Access to a more advanced generative model not only increased task performance but also induced changes in the way that participants wrote prompts, which we highlight through two exploratory analyses. First, prompts to the more advanced model were longer and contained more information. Those assigned to DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and this difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$). Despite this difference in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs. 49% DALL-E 2; $p$ = $0.215$), suggesting that prompts submitted to DALL-E 3 conveyed additional descriptive information. Second, those assigned to DALL-E 3 explored the space of possible prompts differently; they wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses). In other words, those assigned to DALL-E 3 were more likely to thoroughly explore a smaller set of prompting approaches.
The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations must adapt their prompts as models evolve. To quantify how much of the performance improvement came from participants adapting their prompts, we took prompts written by participants in the DALL-E 2 treatment arm and generated images with them using DALL-E 3. This allowed us to isolate the “model effect”—how much better DALL-E 3 performed with prompts written for DALL-E 2 ($\Delta CoSim=0.0084$, $p<10^{-8}$; red in Figure 2), and the “prompting effect”—the portion of the performance ATE due solely to participants adapting their prompting for DALL-E 3 ($\Delta CoSim=0.0079$, $p=0.024$; blue in Figure 2). These effects were nearly equal: 51% and 48% of the performance ATE, respectively. In contrast, we also supplied the prompts written for DALL-E 3 to DALL-E 2. Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$). This implies that DALL-E 3 prompts specifically took advantage of its superior ability to render information, and users naturally discovered these capabilities (Betker et al., 2023).
Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects. The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line). Effects are shown relative to a baseline of prompts written for and replayed on DALL-E 2. Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
Our results show that adapting prompts is crucial for maximizing the benefits of more capable generative models, raising the question of how organizations should approach this adaptation process. One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models. Comparison of our third treatment arm—DALL-E 3 with revision—to the DALL-E 3 arm highlights the potential pitfalls of this approach. We find that automated prompt revisions actually caused participants to perform worse in our context. It reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]). Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). These findings suggest that, as currently implemented, AI-assisted prompt revisions are not a panacea. They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals. Humans, on the other hand, are better at updating their prompts.
[…]
Table 1: Model and Prompting Effect by Performance Decile
Effect Estimate P-value
Model $0.01100$ $<0.00001$
$(0.00186)$
Model $\times$ $-0.00006$ $0.0210$
Performance Decile $(0.00003)$
Prompting $0.01090$ $0.0006$
$(0.00316)$
Prompting $\times$ $-0.00006$ $0.2444$
Performance Decile $(0.00005)$
Discussion
The main limitation of our work is that we only study the transition from DALL-E 2 to DALL-E 3. We leave the study of future transitions and different types of generative AI models to future research. Software systems, models, and prompts are becoming increasingly interwoven, which is a trend we expect to accelerate. If our findings generalize to other models, firms that fail to update their prompts as models improve may forego roughly half of the potential benefits of AI advancement, which will dampen the overall economic impact of AI. Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models. In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities. Such a pattern suggests that technical prompting infrastructure will not be a one-time investment as generative AI models advance. Rather, prompting will be an essential and increasingly intuitive mechanism by which people and firms unlock new models’ capabilities. Fortunately, our findings suggest that humans across the skill distribution can learn to update prompts, even when automated revisions do not suffice.
Material and Methods
Institutional Review. This study was reviewed by the UC Berkeley Committee for the Protection of Human Subjects (CPHS) under Protocol 2023-06-16480.
Outcome Measure. Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image. To compute this, we generated 10 images for each prompt and computed CLIP embeddings (Radford et al., 2021) for both these and the target images. Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
Replay Data. To generate the images used to decompose the prompting and model effects, we regenerated or “replayed” all images with all participant prompts on both models. Note that all main text analyses use replay data from Figure 2—even replays of DALL-E 2 written prompts on DALL-E 2, and DALL-E 3 written prompts on DALL-E 3—for consistency and to avoid model drift issues. All results hold with the “original” data. See SI for details.
Experimental Materials. Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp. Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.
Supplementary Information (SI)

### v5

@@ 2.1 Notation and Problem Setting
so that users cannot exceed perfect fidelity (quality of 1), and each additional unit of effort yields diminishing returns. The user’s utility is then
$U(\theta,s,x)\;=\;Q(\theta,s,x)\;-\;k\,x.$
The user chooses their effort such that it maximizes the utility. We assume $\theta s>k$, which implies there is a unique interior optimum $x^{*}(\theta,s)$
$x^{*}(\theta,s)\;=\;\frac{1}{\theta s}\,\ln\Bigl{(}\frac{\theta s}{k}\Bigr{)}>0,$ (1)
with an optimal quality
$Q^{*}\;=\;Q\big{(}\theta,s,x^{*}(\theta,s)\big{)}=1-\frac{k}{\theta s}>0.$ (2)
Two implications immediately follow from equation 2. Optimal quality is increasing with both model capacity
$\frac{\partial Q^{*}}{\partial\theta}=\frac{k}{\theta^{2}s}>0$ (3)
and user skill
$\frac{\partial Q^{*}}{\partial s}=\frac{k}{\theta s^{2}}>0.$ (4)
Furthermore, the relationship between the changes in these parameters supplies us with a basic prediction, which we can evaluate empirically in our experiment.
Proposition 1.
As model capacity $\theta$ improves, the effect of user skill on optimal quality diminishes.
$\frac{\partial^{2}Q^{*}}{\partial s\partial\theta}=-\frac{k}{\theta^{2}s^{2}}<0$
In other words, improvements in model capacity reduce the performance gap between high and low-skilled users.
2.2 Decomposition into Model and Prompt Effects
As model capacity $\theta$ increases, so does $x^{*}$ in equation 1, implying that a more capable model not only delivers better output for a fixed prompt but also encourages the user to devote more effort to composing prompts. This distinction highlights both a direct capacity-based effect of an improved model, a model effect, and an indirect effect driven by users adapting their prompts to the new model’s potential, a prompting effect.
To see this more formally, suppose the model is upgraded from capacity $\theta_{1}$ to $\theta_{2}$. Let the old prompting effort be $x^{*}(\theta_{1},s)$ and the new (adapted) prompting effort be $x^{*}(\theta_{2},s)$. Define the total improvement in quality $\Delta Q$ is
$\Delta Q\;=\;Q\Bigl{(}(\theta_{2},s,x^{*}(\theta_{2},s)\Bigr{)}\;-\;Q\Bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\Bigr{)}.$
We can decompose $\Delta Q$ into
$\Delta Q$ $={}\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{1},s))\;-\;Q(\theta_{1},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Model Effect }(M)\vphantom{\Big{|}}}$ (5)
$+\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{2},s))\;-\;Q(\theta_{2},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Prompting Effect }(P)\vphantom{\Big{|}}}.$
The first term isolates the gain from simply upgrading the model (holding the user’s prompt strategy fixed at the old optimum), while the second term represents any additional gain from prompt adaptation. In other words, even if $\theta$ improves, failing to adjust how prompts are written could leave substantial performance gains on the table.
This decomposition naturally implies a prediction similar to Proposition 1 for each component of $\Delta Q$ that co-evolve with the changes in model capacity and user skill. In particular, when $k$ is small, the following are proved in Appendix E.1.
Proposition 2.
As user skill $s$ improves, the model effect $M$ decreases and the prompting effect $P$ increases.
Intuitively, when costs are low and model capacity rises, higher-skilled users are already near the upper bound of performance, so the direct model effect they receive may be smaller. In contrast, users who can effectively refine prompts might benefit more from the new capacity, so the prompting effect they receive may be larger. Taken together, these theoretical predictions guide our experimental design, where we randomly vary $\theta$ and measure how users adapt.
3 Experiment Design and Methods
To empirically examine whether users do, in fact, adapt their prompts in response to model improvements—and how much this adaptation contributes to overall performance—we conducted a pre-registered online experiment with 2,059 participants on Prolific between December 12 and December 19, 2023.222Full pre-registration details, including hypotheses and planned analyses, are available in an online repository. Participants were asked to replicate a target image as closely as possible using a generative AI model, with the goal of assessing how both model capability and prompting behavior influence final outcomes. [^3] All procedures were approved by an institutional review board (link removed to preserve author anonymity) and participants provided informed consent. We will release anonymized data and replication code upon publication.
3.1 Experimental Setting
Each participant was randomly and blindly assigned to one of three model conditions: DALL-E 2, DALL-E 3 (hereafter “DALL-E 3 (Verbatim)”), or DALL-E 3 with automatic prompt revision (“DALL-E 3 (Revised)”). These models differ not only in technical capability but also in whether they apply hidden large language model (LLM)-based modifications to user prompts. In the DALL-E 2 condition, participants interacted with an earlier-generation model that interprets prompts directly without intermediate rewriting. In contrast, the DALL-E 3 API, by default, forwards user prompts to GPT-4 before image generation. This intermediate GPT-4 step rewrites the prompt—typically by adding detail or restructuring language—before passing it to the image model. This behavior is intended to improve image quality but occurs silently and without user visibility. In the DALL-E 3 (Revised) condition, we allowed this default behavior to proceed unaltered. In the DALL-E 3 (Verbatim) condition, we attempted to suppress GPT-4-based prompt rewriting by prepending a hidden system message instructing the model to leave the user’s prompt unchanged. While some rewriting still occurred, the rate and extent of modifications were substantially reduced. In all conditions, participants were unaware that their prompt might be rewritten or modified before image generation.
In addition to model assignment, each participant was independently assigned one of 15 target images. These images were drawn from three broad categories—business and marketing, graphic design, and architectural photography—to represent common use cases for text-to-image generation. All images were sourced from platforms that permit free research use (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography) [^4] All of these images have licenses for free use for commercial and noncommercial purposes.. We built a custom interface resembling ChatGPT, with the assigned target image on the right and a scrollable history of prompts and generated images on the left. Participants were explicitly informed that the model was memoryless: every new prompt was processed independently, carrying no information from previous attempts.
Participants had up to 25 minutes to submit prompts, with a requirement to submit at least 10. They were paid \$4 for completing the task, plus a \$8 bonus (a 200% increase) if their highest-scoring image was in the top 20% of participants. The median completion time was 22 minutes, implying an average hourly wage of about \$15. After replicating the target image, participants completed a demographic survey covering age, gender, education, occupation, and self-assessed proficiency in creative writing, programming, and generative AI. We removed from our analyses any participants who did not submit at least 10 prompts, repeated the same prompt five or more times in a row, or failed to complete the post-task survey, resulting in a final sample of 1,893 participants and 18,560 prompts. [^5] Although participants were required to submit at least 10 prompts to be included in our analysis, the final prompt count is slightly below 10 $\times$ 1,893 because we excluded some prompts due to technical issues (e.g., safety filter triggers, duplicate attempt numbers) and limited our main analysis to each participant’s first 10 prompts to mitigate potential selection bias. Full details are provided in Appendix C.4.
3.2 Outcome Definition and Stochastic Generation
The primary outcome in our experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al. 2021). CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together. By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
Because the output of each generative model is stochastic, the same prompt can yield different images across attempts. To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually. We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis. As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments. [^6] DreamSim (Fu et al. 2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods. Our results are robust to using DreamSim in place of CLIP cosine similarity. Full DreamSim-based analyses can be found in Appendix F. The two measures were highly correlated, and our findings were consistent across both.
3.3 Replay Analysis for Separating Model and Prompting Effects
A central goal of our experiment was to distinguish how much of the performance improvement in image replication stems from using a more capable model versus how much comes from users adapting their prompts. Recall that the language of our conceptual framework shows that the total improvement in output quality from upgrading a generative AI model with capacity $\theta_{1}$ to a model with higher capacity $\theta_{2}$ can be written as:
$\Delta Q=Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}-Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)}.$
We decompose this change into two parts. These are the model effect, the gain from applying the same prompts to a better model,
$M=Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}-Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)},$
and the prompting effect, the additional improvement from adapting prompts to take advantage of the more capable model,
$P=Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}-Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}.$
To estimate these components empirically, we conducted an additional analysis using prompts from participants in the DALL-E 2 and DALL-E 3 (Verbatim) conditions. To do so, we took the exact prompts that participants submitted during the experiment and re-submitted (or “replayed”) them to both their originally assigned model and the alternative model, generating new images in each case. [^7] Because the replay was conducted several weeks after the original experiment, all prompts were re-evaluated at the same time, using the same infrastructure, regardless of whether they were being run on the original or alternative model. This ensured that any observed differences in output quality could be cleanly attributed to model capacity and prompt adaptation rather than to changes in model behavior over time (i.e., model drift (Chen et al. 2023)). More specifically, prompts written by DALL-E 2 participants—corresponding to $x^{*}(\theta_{1},s)$—were evaluated both on DALL-E 2 and on DALL-E 3, providing an empirical measure of $Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)}$ and $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$, respectively. This comparison isolates the model effect: the improvement in output quality when keeping prompts fixed and simply upgrading the model. To estimate the prompting effect, we compared the quality of reused DALL-E 2 prompts on DALL-E 3—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$—to the quality of prompts originally written by DALL-E 3 participants and evaluated on the same model—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}$. This captures the additional improvement from users adapting their prompts to better leverage the capabilities of the more advanced system.
4 Results
In this section, we present our empirical findings on how increased model capacity affects user performance in replicating target images. In particular, we focus on three main questions: (i) whether access to a more capable model (DALL-E 3) improves performance; (ii) how users modify their prompts in response to the improved model; (iii) and how much of the overall performance gain can be attributed to model improvements versus prompt adaptation. Finally, we compare these empirical results to the predictions of our conceptual framework. Our analysis draws on both the main experimental data and the replay data described above.
4.1 Overall Impact of Model Upgrades
We begin by examining whether participants using DALL-E 3 achieve higher performance than those using DALL-E 2 as implied by Equation 3. Figure 1 summarizes these findings. Panel A presents three representative target images and, for each, three generated images drawn from the full sample of participants across both model conditions. The middle row for each target shows the image whose cosine similarity to the target is closest to the mean similarity across all participants. The rows above (below) show images that are approximately one average treatment effect (ATE) more (less) similar to the target than the mean. This visualization provides qualitative intuition for the magnitude of the effect we estimate: the typical difference in fidelity between participants using DALL-E 2 and those using DALL-E 3 (Verbatim). Panel B shows that, across the 10 required prompt attempts, participants assigned to DALL-E 3 (Verbatim) produce images that are, on average, 0.0164 higher in cosine similarity to the target (95% CI: [0.0104, 0.0224], $p<10^{-5}$). This improvement corresponds to roughly 0.19 standard deviations in performance. The gap persists across all attempts; participants using DALL-E 3 start off producing closer matches and maintain that edge through their 10th prompt.
Figure 1: Overall Performance and Prompting Behavior. (A) For three example target images, the middle row shows participant-generated images closest to the mean similarity across all prompts. The rows above and below show images of approximately one average treatment effect (ATE) more or less similar to the target, illustrating the typical performance difference between model conditions. (B) Top: average CLIP cosine similarity to the target image by attempt, separately for DALL-E 2 and DALL-E 3 participants. Bottom: the difference between these averages, with the dark blue line indicating the overall ATE and the shaded region showing the 95% confidence interval. (C) Average prompt length by attempt (y-axis), with error bars representing 95% confidence intervals. Color shading indicates the average textual similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
Participants’ dynamic prompting behavior also differs substantially between the two models. As shown in panel C of Figure 1, those assigned to DALL-E 3 write prompts that are, on average, 24% longer than those assigned to DALL-E 2, and this gap widens over successive attempts. Moreover, we observe that DALL-E 3 participants are more likely to reuse or refine their previous prompts (indicated by the color scale), which suggests a more exploitative approach once they discover the model’s capacity to handle detailed or complex instructions. Analyses of parts of speech confirm that these extra words likely provide additional descriptive information rather than mere filler: the proportion of nouns and adjectives—the two most descriptively informative parts of speech—is nearly identical across model conditions (48% for DALL-E 3 vs. 49% for DALL-E 2; $p=0.215$). This suggests that the increase in prompt length reflects the addition of semantically rich content rather than unnecessary verbosity.
4.2 Replay Analysis and Decomposition of Effects
The differences we observe in prompting behavior suggest that users are actively adapting to the capabilities of the model they are assigned. But how much of the overall performance improvement we observe for DALL-E 3 users is due to the model’s enhanced technical capacity, and how much is due to users rewriting their prompts in response to that capacity? To answer this question, we turn to the replay analysis described earlier, which allows us to isolate these two effects identified in Equation 5 empirically.
Panel A of Figure 2 presents the results. To estimate the model effect, we compare the performance of prompts originally written by DALL-E 2 participants when evaluated on DALL-E 2 (the model they were written for) versus when evaluated on DALL-E 3 (Verbatim). Because these prompts were written without knowledge of DALL-E 3’s capabilities, any improvement reflects the gain from using a more capable model while holding the prompt fixed. We find that performance improves by 0.0084 in cosine similarity when these prompts are evaluated on DALL-E 3 ($p<10^{-8}$; bootstrapped standard errors clustered at the participant level), which accounts for approximately 51% of the total difference in performance between the DALL-E 2 and DALL-E 3 arms.
To estimate the prompting effect, we then compare the performance of these same DALL-E 2 prompts to the performance of prompts originally written by DALL-E 3 participants, both evaluated on DALL-E 3. Because both sets of prompts are evaluated on the same model, any difference reflects the effect of users adapting their prompts to the model’s capabilities. We find that this prompting effect accounts for the remaining 48% of the total improvement, corresponding to an increase of 0.0079 in cosine similarity ($p=0.024$). The total treatment effect—i.e., the difference between the original DALL-E 2 prompts on DALL-E 2 and the original DALL-E 3 prompts on DALL-E 3—is 0.0164.
Importantly, when we apply prompts written by DALL-E 3 users to DALL-E 2, we observe no performance benefit relative to the original DALL-E 2 prompts ($\Delta=0.0020$; $p=0.56$). This asymmetry reinforces the idea that the gains from prompt adaptation depend on the model’s capacity to act on that additional information.
Figure 2: Replay Analysis and Effect Decomposition. (a) Average performance of prompts evaluated across four prompt-model combinations. Comparing DALL-E 2 prompts evaluated on DALL-E 2 versus DALL-E 3 isolates the model effect. Comparing reused DALL-E 2 prompts on DALL-E 3 to original DALL-E 3 prompts on DALL-E 3 isolates the prompting effect. Error bars represent 95% confidence intervals based on bootstrapped standard errors clustered at the participant level. (b) A single target image with two submitted prompts: one written by a DALL-E 2 participant (top row) and one by a DALL-E 3 participant (bottom row). Images show how each prompt performs on both models, illustrating the model and prompting effects qualitatively.
Panel B of Figure 2 illustrates these effects using a single target image. The two rows show different prompts submitted for that target, along with the images they generate when evaluated on each model. In the top row, a prompt originally written by a DALL-E 2 participant yields a higher-fidelity image when replayed on DALL-E 3, demonstrating the improvement in output quality that comes from upgrading the model while holding the prompt fixed. In the bottom row, a prompt written by a DALL-E 3 participant produces a noticeably lower-quality image when rendered by DALL-E 2, underscoring the limits of prompt adaptation when the model lacks sufficient capacity to execute the instructions effectively.
Taken together, these findings offer empirical support for our theoretical claim: prompt adaptation operates as a dynamic complement that users deploy in response to improved model capabilities—and accounts for a substantial share of realized performance gains.
4.3 Skill Heterogeneity
Effect Estimate (SE) p-value
Total 0.0224 (0.00312) $<$ 0.00001
Total $\times$ Performance Decile -0.000115 (0.000047) 0.0152
Model 0.0113 (0.00186) $<$ 0.00001
Model $\times$ Performance Decile -0.000059 (0.000025) 0.0210
Prompting 0.0109 (0.00316) 0.000564
Prompting $\times$ Performance Decile -0.000056 (0.000048) 0.244
Table 1: Regression estimates for the total, model and prompting effects, and their interactions with participant performance decile. The total and model effects are significantly larger for lower-skilled participants, as indicated by the negative and statistically significant interaction term. The prompting effect does not vary significantly across deciles, suggesting that participants across the skill distribution adapt their prompts similarly.
Table 1 presents the results of a regression analysis that tests whether the model, prompting and total effects vary systematically across participants of differing skill levels. We regress performance on indicators for model, prompting and total effects, as well as their interactions with participants’ performance decile (ranging from lowest to highest).
We make three observations. First, the interaction between total effect and performance decile is negative and statistically significant ($-0.000115$, $p=0.0152$). This suggests that model improvements reduce the overall gap between high and low-performing users, consistent with proposition 1 in our conceptual framework. Second, the interaction between the model effect and performance decile is also negative and statistically significant ($-0.000059$, $p=0.0210$). This suggests that the model effect mainly benefits the lower-performing users—consistent with our theoretical predictions in proposition 2 and due to diminishing returns for those already near the performance ceiling. Third, we do not find any evidence that the benefits of prompt adaptation vary across the skill distribution in a meaningful way since the interaction between the prompting effect with performance decile is not statistically significant ($-0.000056$, $p=0.2444$). While this contrasts with our theoretical prediction in Proposition 2—which anticipated that higher-skill participants would benefit more from prompt adaptation—it is possible that the effect exists but is modest in size and difficult to detect given the statistical power of our design.
4.4 Prompt Revision
We also evaluate whether an automated prompt adaptation system can substitute for human-driven adaptation. In the DALL-E 3 (Revised) condition, user prompts were silently rewritten by GPT-4 before being submitted to the image model. Although participants in this condition still outperformed those using DALL-E 2 ($\Delta=0.0069$; $p=0.042$), they achieved substantially lower performance than participants using DALL-E 3 without revision. On average, automated prompt revision reduced the benefit of DALL-E 3 by 58% (95% CI: [40%, 76%]). Manual inspection of the revised prompts in the DALL-E 3 (Revised) arm confirms that GPT-4 often added extraneous details or subtly altered the intended meaning of participant instructions, leading to degraded performance. In this case, the automatic rewriting process was not well-aligned with the user’s goal of replicating a specific target image as faithfully as possible. More broadly, this result highlights a potential risk of using invisible system-level prompt rewriting: when the objectives of the rewriting mechanism are not tightly aligned with user intent, even well-intentioned modifications can interfere with the user’s ability to effectively guide the model.
Together, these findings demonstrate that the performance gains observed when users are given access to a more capable generative model stem from both improvements in model architecture and changes in user behavior. Participants assigned to DALL-E 3 not only achieve higher performance but also write longer, more refined prompts over time. Our replay analysis shows that these adapted prompts account for nearly half of the overall performance improvement, with the remainder attributable to the model’s enhanced rendering capabilities. This pattern holds across the skill distribution: while lower-skilled users benefit more from model upgrades, prompt adaptation contributes consistently across all levels of baseline performance. Finally, automated prompt rewriting—when misaligned with user intent—fails to replicate the benefits of direct user adaptation. These results highlight the central role that prompting behavior plays in shaping how users interact with and benefit from improved generative AI systems.
5 Discussion and Conclusion
Our findings contribute to the literature on human–AI collaboration and the economics of technology by highlighting the role of prompt adaptation as a dynamic complement that co-evolves with model improvements (Böhm and Schedlberger 2023). As generative AI models advance—often substantially from month to month—organizations that fail to adapt their prompting strategies may forgo a meaningful share of the economic value these upgrades make possible.
This co-evolutionary pattern is not unique to AI. It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (Brynjolfsson 1993, David 1990). However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter. Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (Brynjolfsson and Hitt 2000, Brynjolfsson et al. 2021). Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—accounts for roughly half of the realized performance gains associated with a model upgrade.
This finding extends research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997) and post-adoptive IT use, which emphasizes how value emerges through user experimentation and learning over time (Jasperson et al. 2005). Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill. Participants in our study—none of whom were expert prompt engineers—improved performance through trial-and-error within a single session. This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (Attewell 1992, Von Hippel 2006). The accessibility of prompt adaptation suggests a path toward more broadly distributed productivity gains—provided that users are supported by scaffolds and interfaces that enable iterative refinement (Rogers 2003).
At the same time, this accessibility introduces risk. Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve. This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (Henderson and Clark 1990). Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
Finally, our results caution against automated prompting systems that intervene without aligning with user intent. In our experiment, DALL-E 3’s automatic prompt rewriting—designed to enhance usability—reduced realized performance gains by 58%. This does not imply that automated prompting is inherently counterproductive. In some settings, automation may simplify use or improve consistency. But when the goals of a rewriting system diverge from those of the user—as in our task, where general improvements (e.g., aesthetic quality or detail) were prioritized over faithful image replication—such interventions can interfere with the adaptive strategies users are developing. Similar challenges have been observed in other domains: for instance, in clinical documentation, automated prompt optimization improved consistency but still required manual revision to ensure quality and alignment (Yao et al. 2024). These findings reinforce the broader insight—rooted in both human–AI collaboration and dynamic capabilities research—that adaptive user behavior is not merely a response to system performance but a critical input. Interventions that disrupt this adaptive process risk eroding the very dynamic complements that make advanced systems productive (Fügener et al. 2022).
While this study offers new insight into prompt adaptation, it has several limitations. First, our focus was restricted to a single transition (from DALL-E 2 to DALL-E 3) and one type of generative AI (text-to-image generation). Although our conceptual framework suggests these mechanisms may generalize to other domains, further research is needed to assess how these dynamics play out in text generation, programming, scientific research, and other high-stakes settings. Second, we observe short-run adaptation behavior in a controlled setting, whereas longer-term learning dynamics, organizational feedback structures, or team-based workflows may shape prompting strategies differently in real-world environments. Third, while our replay analysis isolates the effect of prompt adaptation, we do not identify which specific types of prompt modifications (e.g., lengthening, lexical substitution, structural rephrasing) have a causal effect on performance. Future work should investigate these finer-grained mechanisms to better inform prompting practices and interface design.
Building on these limitations, several promising research directions emerge. One direction is to examine prompt adaptation longitudinally—tracking how users develop durable heuristics that generalize across domains and respond to shifting model behavior over time. Another line of inquiry involves organizational-level complements to prompting, such as shared prompt repositories, collaborative refinement practices, or analytics dashboards that surface effective patterns. A third area concerns interface design—specifically, how features like auto-complete, prompt scoring, or real-time feedback influence users’ ability to experiment and adapt. Finally, researchers might explore how organizations balance standardization (which streamlines processes) with adaptability in prompting workflows, in order to avoid creating rigid routines that lag behind technical advances.
For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance. The near-equal contributions of model improvements and prompt adaptation suggest that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (Arthur 2009, Dosi 1982). As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.